OpenStudy (jkbo):

What is the slope-intercept form of the equation y – 7 = – 5/2 (x + 4)? a.y – 2 = –5/2(x + 2) b.y + 7 = –x + 5/2 c.y = –4/7+ 2 d.y = –5/2x – 3

OpenStudy (anonymous):

y = mx + b Simplify what you have and leave the y isolated. y - 7 = -5/2(x + 4) y - 7 = -5/2x - 10 y = ?

OpenStudy (jkbo):

i dont know how to simplify this